# Shortest-path percolation on scale-free networks

### v1

@@ Acknowledgements.
We acknowledge the support of the AccelNet-MultiNet program, a project of the National Science Foundation (Award No. 1927425 and No. 1927418). This work received partial support from the Air Force Office of Scientific Research (Grant No. FA9550-24-1-0039). CC acknowledges the PRIN project No. 20223W2JKJ “WECARE”, CUP B53D23003880006, financed by the Italian Ministry of University and Research (MUR), Piano Nazionale Di Ripresa e Resilienza (PNRR), Missione 4 “Istruzione e Ricerca” - Componente C2 Investimento 1.1, funded by the European Union - NextGenerationEU. The funders had no role in study design, data collection and analysis, the decision to publish, or any opinions, findings, conclusions, or recommendations expressed in the manuscript.
@@
Figure S4: Pseudo-critical point $r_{c}(N)$. We plot the pseudo-critical point $r_{c}(N)$ as a function of network size $N$. Results are valid for SFNs with (a) $\lambda=2.1$, (b) $\lambda=2.7$, (c) $\lambda=3.5$, and (d) $\lambda=4.5$. Note that the horizontal dashed lines in panel (c) and (d) are the plateau values of $r_{c}(N)$ for SPP with $C=N$. These values are obtained by taking the average of $r_{c}(N)$ with $N>10^{5}$, which are within the error bounds reported in Table 1.
Figure S5: Data Collapse of probability distributions of critical observables. Similar to Fig. 4 in the main, but results are valid for SPP with $C=1$. Note that this corresponds to ordinary bond percolation.
Figure S6: Data Collapse of probability distributions of critical observables. Similar to Fig. 4 in the main, but results are valid for SPP with $C=3$.
Figure S7: Data Collapse of probability distributions of critical observables. Similar to Fig. 4 in the main, but results are valid for SPP with $C=N$.
Figure S8: Data Collapse of $P$ and $S$. We plot the rescaled observable $PN^{\beta/\bar{\nu}}$ as a function of (a) $(r-\hat{r}_{c})N^{1/\bar{\nu}_{1}}$ and (b) $(r-\hat{r}_{c})N^{1/\bar{\nu}_{2}}$, respectively. Similarly, we plot the rescaled observable $SN^{-\gamma/\bar{\nu}}$ as a function of (c) $(r-\hat{r}_{c})N^{1/\bar{\nu}_{1}}$ and (d) $(r-\hat{r}_{c})N^{1/\bar{\nu}_{2}}$, respectively. Results are valid for SFNs with $\lambda=2.1$ and $k_{min}=4$, and SPP model with $C=2$. Note that in order to make a fair comparison between the data collapse using $1/\bar{\nu}_{1}$ and $1/\bar{\nu}_{2}$, the range of the abscissa is adjusted to display 200 bins of the curve for $N=2^{20}$.

### v2

@@ Acknowledgements.
We acknowledge the support of the AccelNet-MultiNet program, a project of the National Science Foundation (Awards No. 1927425 and No. 1927418). This work received partial support from the Air Force Office of Scientific Research (Grant No. FA9550-24-1-0039). CC acknowledges the PRIN Project No. 20223W2JKJ “WECARE”, CUP B53D23003880006, financed by the Italian Ministry of University and Research (MUR), Piano Nazionale Di Ripresa e Resilienza (PNRR), Missione 4 “Istruzione e Ricerca” - Componente C2 Investimento 1.1, funded by the European Union - NextGenerationEU. The funders had no role in study design, data collection and analysis, the decision to publish, or any opinions, findings, conclusions, or recommendations expressed in the manuscript.
Data Availability
The source code that supports the findings of this article is openly available [39],
Appendix A Complementary analysis for Sec. IV.2
To provide additional evidence that the SPP process effectively homogenizes the heterogeneous network structure before the percolation transition takes place, we perform the following analysis. Consider the following quantity:
$\rho(N,\ r)=\frac{\langle k^{2}\rangle}{\langle k\rangle^{2}},$ (15)
where $N$ is the size of the network, $\langle k\rangle$ is the average degree after removing a fraction $r$ of the edges, and $\langle k^{2}\rangle$ is the average of the squared degrees under the same removal. Note that we omitted the $r$-dependence of $k$ to avoid overcrowding the equation. $\rho(N,\ r)$ effectively captures the degree heterogeneity of the network.
Now, we compare $\rho(N,\ r)$ of the initial network, $\rho^{o}(N)=\rho(N,\ r=0)$, with the value obtained after removing a fraction $r=r^{\dagger}$ of edges before the percolation transition takes place, $\rho^{\dagger}(N)=\rho(N,\ r=r^{\dagger})$. Then, we define the quantity $\Omega(N)$ and assume the functional form as follows:
$\bigg\langle\frac{\rho^{\dagger}(N)}{\rho^{o}(N)}\bigg\rangle\equiv\Omega(N)=\Omega_{o}+aN^{-\xi},$ (16)
similar to Eq. (14). To properly complement the analysis, we choose $r=r^{\dagger}$ (i) at a fixed value of $r^{\dagger}=q$ or (ii) $r^{\dagger}=qr_{c}^{*}(N)$ for each realization, using q=0.7, 0.8,q=0.7,\ 0.8, and $0.9$. We also add a subscript to distinguish the quantities in Eq. (16), i.e., $\Omega_{1}(N)$ and $\Omega_{2}(N)$. Note that nodes with zero degree are excluded from this computation.
We perform the analysis on SFNs with $\lambda=2.1$. In Fig. 6, we plot $\Omega(N)$ as a function of $N$ for different values of $q$ and $C$. We also report the estimated variables $\Omega_{o,1}$, $\Omega_{o,2}$, $\xi_{1}$, and $\xi_{2}$ in Table S4 of Ref. [31]. These results clearly complement the analysis in Fig. 6, indicating that the SPP process with $C>1$ effectively homogenizes the heterogeneous structure of SFNs.
Figure 6: Scaling of the $\rho$ ratios. Plot of $\Omega_{1}(N)$ as a function of $N$ for different values of $C$. Results are valid for SFNs with $\lambda=2.1$ using (a) $q=0.7$ (c) $q=0.8$, and (e) $q=0.9$. We also plot $\Omega_{2}(N)$ as a function of $N$ for different values of $C$ using (b) $q=0.7$ (d) $q=0.8$, and (f) $q=0.9$. Plateau values and exponents are reported in Table S4 of Ref. [31]. Symbols are the average values and error bars are the standard deviations. For the scaling analysis, we use 50 realizations for each data point. We also drop the data point corresponding to the smallest network size when performing the scaling analysis.
@@
Figure S4: Pseudocritical point $r_{c}(N)$. We plot the pseudocritical point $r_{c}(N)$ as a function of network size $N$. Results are valid for SFNs with (a) $\lambda=2.1$, (b) $\lambda=2.7$, (c) $\lambda=3.5$, and (d) $\lambda=4.5$. Note that the horizontal dashed lines in panel (c) and (d) are the plateau values of $r_{c}(N)$ for SPP with $C=N$. These values are obtained by taking the average of $r_{c}(N)$ with $N>10^{5}$, which are within the error bounds reported in Table 1.
Figure S5: Estimation of $1/\bar{\nu}$. Similar to Fig. 4 in the main, but results are valid for SPP with $C=1$. Note that this corresponds to ordinary bond percolation.
Figure S6: Estimation of $1/\bar{\nu}$. Similar to Fig. 4 in the main, but results are valid for SPP with $C=3$.
Figure S7: Estimation of $1/\bar{\nu}$. Similar to Fig. 4 in the main, but results are valid for SPP with $C=N$.
Figure S8: Data Collapse of $P$ and $S$. We plot the rescaled observable $PN^{\beta/\bar{\nu}}$ as a function of (a) $r-r_{c}(N)]N^{1/\bar{\nu}_{1}}$ and (b) $r-r_{c}(N)]N^{1/\bar{\nu}_{2}}$, respectively. Similarly, we plot the rescaled observable $SN^{-\gamma/\bar{\nu}}$ as a function of (c) $r-r_{c}]N^{1/\bar{\nu}_{1}}$ and (d) $r-r_{c}(N)]N^{1/\bar{\nu}_{2}}$, respectively. Results are valid for SFNs with $\lambda=2.1$ and $k_{min}=4$, and SPP model with $C=2$. Note that in order to make a fair comparison between the data collapse using $1/\bar{\nu}_{1}$ and $1/\bar{\nu}_{2}$, the range of the abscissa is adjusted to display 200 bins of the curve for $N=2^{20}$.
@@ 3 $2^{10}$$2^{17}$ 300
$q$ $C$ $\Omega_{o,1}$ $\xi_{1}$ $\Omega_{o,2}$ $\xi_{2}$
0.7 1 0.889(1) 0.330(3) 0.884(1) 0.287(2)
2 0.000(1) 0.209(4) 0.000(1) 0.198(5)
3 0.000(1) 0.210(3) 0.000(1) 0.198(3)
NN 0.000(1) 0.190(2) 0.000(1) 0.178(3)
0.8 1 0.779(1) 0.327(2) 0.768(1) 0.275(2)
2 0.000(1) 0.235(5) 0.000(1) 0.224(5)
3 0.000(1) 0.236(4) 0.000(1) 0.223(4)
NN 0.000(1) 0.220(4) 0.000(1) 0.206(4)
0.9 1 0.560(1) 0.308(1) 0.535(1) 0.252(1)
2 0.000(1) 0.251(6) 0.000(1) 0.247(6)
3 0.000(1) 0.251(6) 0.000(1) 0.246(5)
NN 0.000(1) 0.244(6) 0.000(1) 0.236(5)
Table S4: Scaling of the $\rho$ ratios. We report the estimated variables $\Omega_{o,1}$, $\Omega_{o,2}$, $\xi_{1}$, and $\xi_{2}$ using Eq. (16).